# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v4

@@ IV-C2 Evaluating other components
Based on all ablation studies, it demonstrates that our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance issue of each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection. The introduction of each component proves beneficial for performance improvement, with NS emerging as the most effective component. Moreover, the results on each AU are also shown in Table VII. It can be seen that the results on basically every AU have improved or approached after we add , PS and NS.
IV-D Hyperparameters Analysis
@@
IV-D1 , the probabilities of various positive sample pairs
@@
The experimental results are presented in Table VIII. When =0.15, =0.15, =0.7, and =0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The conclusion is consistent with [89, 90]. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
IV-D2 Backpropagation rate controller
@@
V Conclusion
In this paper, we introduce a contrastive learning framework for AU detection. Our proposed contrastive loss, named AUNCE, shares similarities with InfoNCE, transforming the task of learning entire pixel-level information of AUs into the more effective learning of subtle feature differences when AUs activate. Additionally, we present a positive sample sampling strategy that introduces three types of positive sample pairs, addressing the challenge of AU labels tainted by noise and errors in two AU datasets. To further enhance the framework, we incorporate a negative sample re-weighting strategy specifically aimed for minority AUs. This strategy controls the relative contribution of minority class samples to the parameter updates, effectively addressing the class imbalance issue of each AU type, and contributing to hard negative sample mining. Experimental results underscore the superior performance of AUNCE in comparison to existing methods.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Science and Technology Program (Nos. 2024NSFTD0036, 2024ZHCG0166). The computation is completed in the HPC Platform of Huazhong University of Science and Technology.

### v5

@@ 4.3.2 Evaluating other components
Based on all ablation studies, it demonstrates that our proposed discriminative contrastive learning paradigm enhances the encoding capability of the shared encoder, improves classification accuracy by addressing the class imbalance issue of each AU type, and bolsters model robustness against AU labels affected by noise and errors. This, in turn, leads to enhanced generalization performance in AU detection. The introduction of each component proves beneficial for performance improvement, with NS emerging as the most effective component. Moreover, the results on each AU are also shown in Table 8. It can be seen that the results on basically every AU have improved or approached after we add , PS and NS.
4.4 Hyperparameters Analysis
@@
4.4.1 , the probabilities of various positive sample pairs
@@
The experimental results are presented in Table 9. When =0.15, =0.15, =0.7, and =0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The conclusion is consistent with [55]. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.
4.4.2 Backpropagation rate controller
@@
5 Conclusion
In this paper, we proposes a discriminative contrastive learning framework tailored for AU detection, introducing a novel contrastive loss named AUNCE. AUNCE builds upon the InfoNCE loss, reformulating the task from learning pixel-level information of the entire face to focusing on the subtle feature variations that occur during AU activation, thereby enhancing detection efficiency. We further introduce a positive sample sampling strategy that constructs three types of positive sample pairs, effectively addressing the challenge of AU labels tainted by noise and errors in four AU datasets. To combat the class imbalance problem often seen in AU detection, we implement a negative sample re-weighting strategy specifically targeting minority AUs, facilitating the identification of hard negative samples. Experimental results reveal that AUNCE consistently outperforms existing methods, particularly those employing weighted cross-entropy loss. However, our negative sample reweighting strategy may inadvertently hinder the model’s ability to learn features of common AUs. An excessive emphasis on minority classes could result in underfitting of the majority classes. To overcome this limitation, our future research will focus on developing dynamic reweighting techniques with adaptive mechanisms that balance minority and majority classes, enhancing the robustness and adaptability of our framework across diverse conditions.
@@ Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62176221, 61572407), and Sichuan Province Science and Technology Support Program (Nos. 2024NSFTD0036, 2024ZHCG0166). The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used in order to polish our paper. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.